OpenStudy (anonymous):

Help please!!! A rectangular garden is 8m long and 4m wide. When the garden is widened on all four sides by x meters, the area is triples. Find the x

hartnn (hartnn):

original Area=A =8*4 new dimensions : (8+2x) and (4+2x) new area = (8+2x)(4+2x) = 3A = 8*4 so, can u solve for x in (8+2x)(4+2x) = 8*4

hartnn (hartnn):